State and prove the weierstrass M -test for the uniform convergence of a series of functions


1
Expert's answer
2021-03-25T13:19:33-0400

Statement: A series "\sum^{\infin}_1U_n(x)" of functions will converge uniformly on I. If there exist a convergent series "\sum^{\infin}_1M_n" of positive constant such that "|U_n(x)|\eqslantless M_n \forall n \in N and \space \forall \in I"


Proof: "\sum M_n" is convergent

By Cauchy's principle for given "\in >0 \& m \in N" such that "M_{n+1}+M_{n+2}+.....+M_{n+p} < \in ..........1"

"\forall n \eqslantgtr m; p=1,2,3"

Our hypothesis "|U_n(x)|\eqslantless M_n \forall n \in N and \space \forall \in I"

Now "|U_{n+1}(x)+U_{n+2}(x)+...+U_{n+p}(x)| \eqslantless |U_{n+1}(x)|+|U_{n+2}(x)|+...+|U_{n+p}(x)| \eqslantless M_{n+1}+M_{n+2}....+M_{n+p} < \in \forall n \eqslantgtr m , p=1,2..."

"|U_{n+1}(x)+U_{n+2}(x)+...+U_{n+p}(x)| < \in \forall n < m , p=1,2..."

Therefore "U_n(x)" convergent uniformly and absolutely on I